# Analyzing GitHub Issues and Pull Requests in nf-core Pipelines: Insights into nf-core Pipeline Repositories

### v1

@@ Abstract.
Scientific Workflow Management Systems (SWfMSs) such as Nextflow have become essential software frameworks for conducting reproducible, scalable, and portable computational analyses in data-intensive fields like genomics, transcriptomics, and proteomics. Building on Nextflow, the nf-core community curates standardized, peer-reviewed pipelines that follow strict testing, documentation, and governance guidelines. Despite its broad adoption, little is known about the challenges users face during the development and maintenance of these pipelines. This paper presents an empirical study of 25,173 issues and pull requests from these pipelines to uncover recurring challenges, management practices, and perceived difficulties. Using BERTopic modeling, we identify 13 key challenges, including pipeline development and integration, bug fixing, integrating genomic data, managing CI configurations, and handling version updates. We then examine issue resolution dynamics, showing that 89.38% of issues and pull requests are eventually closed, with half resolved within three days. Statistical analysis reveals that the presence of labels (large effect, $\delta$ = 0.94) and code snippets (medium effect, $\delta$ = 0.50) significantly improve resolution likelihood. Further analysis reveals that tool development and repository maintenance poses the most significant challenges, followed by testing pipelines and CI configurations, and debugging containerized pipelines. Overall, this study provides actionable insights into the collaborative development and maintenance of nf-core pipelines, highlighting opportunities to enhance their usability, sustainability, and reproducibility.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The 29th International Conference on Mining Software Repositories; 13–14 April, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Scientific Workflow Management Systemsccs: Applied computing Bioinformatics Pipelines
@@ 1. Introduction
The rapid growth of high-throughput sequencing and other omics (e.g., genomics, transcriptomics metagenomics) technologies has led to an explosion of large, heterogeneous, and complex bioinformatics datasets (Yamada et al., 2021). Analyzing these data requires chaining together multiple computational tools (modules) with diverse dependencies, formats, and parameter configurations. To address these challenges, scientific workflow management systems (SWfMSs) such as Nextflow (Di Tommaso et al., 2017), Snakemake (Köster and Rahmann, 2012), and Galaxy (53) have emerged as indispensable software frameworks, enabling researchers to design, execute, and share multi-step computational analyses with improved reproducibility, scalability, and portability (Deelman et al., 2009; Ewels et al., 2020a).
One of the most widely adopted SWfMSs is Nextflow (Di Tommaso et al., 2017), which provides a domain-specific language for defining pipelines and integrates seamlessly with package managers (e.g., Conda) and containerization technologies (e.g., Docker, Singularity). Nextflow also supports deployment on various infrastructures, including local workstations, high-performance computing (HPC) clusters, and cloud platforms. This flexibility allows developers to prototype pipelines locally and subsequently scale them to large computational resources with minimal configuration changes.
@@
From a development standpoint, nf-core offers a standardized pipeline template that enforces coding style and guidelines, reducing entry barriers and promoting code consistency. Automated synchronization propagates template updates, while governance, code reviews, and integration with Bioconda/conda-forge support collaborative and sustainable development. Despite these strengths, maintaining nf-core pipelines remains challenging. Contributors must manage dependencies across diverse environments, resolve synchronization conflicts, update documentation, and ensure rigorous testing. These tasks are coordinated primarily through GitHub issues and pull requests (PRs), which facilitate bug reports, feature requests, and discussions. While prior studies have examined collaboration and maintenance through GitHub artifacts (Yang et al., 2023; Li et al., 2021; Alam et al., 2025), systematic analyses of the types of questions and issues raised within nf-core’s bioinformatics workflows are still lacking.
@@
In this paper, we address this gap through an empirical study of GitHub issues and PRs from nf-core pipeline repositories. Using BERTopic modeling, we identify 13 key challenge areas, including pipeline development and integration, bug fixing, genome data integration, CI management, and version updates, and examine how developers manage and resolve them. Our analysis shows that 89.38% of issues and PRs are resolved, with 50.74% closed within three days. Across 125 active repositories (12 are archived out of 137 total as of August 22, 2025), three have no open issues or PRs, and two have none closed. About 4.64% of issues and PRs remain entirely ignored. We also observe that 92.74% (115) of repositories utilize labeling, while 90.32% (112) assign issues and PRs to specific contributors. Beyond the default labels, we find three additional frequently used labels among the top ten, namely WIP, DSL2, and docs. Among closed cases, 60.88% are self-closed, whereas only 17.13% involve assigned contributors. Furthermore, our analysis indicates that issues and PRs with labels and the inclusion of code blocks are more likely to be closed, suggesting that providing proper labels and essential code snippets, can enhance the clarity and resolution of reported issues and PRs. We make the following contributions in this paper:
@@
  • We conduct an empirical study to systematically investigate issues and PRs in open-source analysis pipelines within the nf-core ecosystem. By analyzing 25,173 entries from 125 nf-core pipeline repositories, our study provides the community with a comprehensive dataset that facilitates a deeper understanding of collaboration, maintenance, and development practices in large-scale scientific software projects.
@@
  • We categorize user-reported challenges into a taxonomy of 13 topics, offering insights into recurring problems and development practices. This categorization aims to support developers in better understanding user challenges, thereby guiding improvements in pipeline usability, functionality, and overall maintenance.
@@
  • Our findings highlight the value of good practices in managing issues and PRs. We recommend active use of GitHub features (e.g., labeling, assigning) for organization and accountability, and encourage contributors to provide clear, well-formatted descriptions with code snippets to improve clarity and resolution efficiency.
@@ 2. Background and Related Work
2.1. SWfMSs and Pipelines
A SWfMS is a specialized software platform that enables researchers to design, execute, and manage complex computational experiments often referred to as scientific workflows (Liu et al., 2015). These systems automate repetitive tasks, integrate diverse tools, and ensure reproducibility by tracking data provenance. Examples include Galaxy (53), Nextflow (Di Tommaso et al., 2017), Snakemake (Köster and Rahmann, 2012), Pegasus (Deelman et al., 2015), and Taverna (Oinn et al., 2004).
A pipeline is a concrete instance of a scientific workflow (Barker and Van Hemert, 2007), representing a sequence of data processing steps tailored to a specific research task, such as RNA-seq analysis in bioinformatics. Pipelines are typically written in workflow languages like domain specific language (DSL) (Fowler, 2010), common workflow language (CWL) (Amstutz et al., 2016), workflow description language (WDL), or yet another workflow language (YAWL) (Van Der Aalst and Ter Hofstede, 2005) and are often modular and reusable. Community repositories such as nf-core curate standardized pipelines. Examples include rnaseq (Ewels et al., 2020b), sarek (Hanssen et al., 2024) and methylseq (nf-core methylseq authors, 2024). In short, the SWfMS is the platform, while the pipeline is the recipe. For example, Nextflow provides the engine to run RNA-seq pipeline, ensuring scalability, reproducibility, and efficient execution.
@@ 2.1. SWfMSs and Pipelines
The nf-core project develops reproducible bioinformatics pipelines using Nextflow (Ewels et al., 2020a). Supported by many leading organizations such as Seqera, SciLifeLab, the Center for Genomic Regulation, AWS, and Microsoft Azure, it ensures high-quality, portable analyses across diverse environments. Through shared maintenance, peer review, and centralized management of reusable modules and subworkflows, it fosters collaboration and enhances reuse. As of August 22, 2025, it hosts 125 active pipelines and over 1,600 modules (nf-core community, 2025a), exemplifying a sustainable, community-led model for bioinformatics research (nf-core community, 2025b).
@@ 2.3. Tracking Issues and Pull Requests in GitHub
Software systems are not always reliable (Gundersen et al., 2018), and maintaining open-source repositories presents ongoing challenges. To facilitate collaboration, GitHub introduced Issues in 2009 (Preston-Werner, 2009) for reporting bugs, discussing features, and asking questions, and Pull Requests in 2008 (GitHub Marketplace, 2025) to propose, review, and merge code changes. Figure 1 illustrates an issue from the nf-core/mag [^2] https://github.com/nf-core/mag/issues/378 pipeline designed for the assembly, binning, and annotation of metagenomes. In this example, the individual who created the issue, referred to as the issue raiser, provides a title and a description that specify the reported concern. Issues can exist in either an open or closed state, reflecting whether they remain unresolved or have been addressed. Any GitHub user with access to the repository may contribute to the discussion through comments, which often provide clarifications, additional details, or potential solutions. For management purposes, issues are supported by two fundamental mechanisms: labels and assignees. Labels can be created by contributors with write access to the repository and applied or removed by those with triage permissions, thereby enabling classification and prioritization. Assignees, by contrast, can be designated by users with write access, ensuring that responsibility for resolution is explicitly assigned. All actors involved in the lifecycle of an issue, including the raiser, commenters, and assignees, are collectively regarded as participants. Both issues and PRs share features such as titles, descriptions, labels, and threaded discussions. PRs, however, add code-specific functionality, branch linking, commit and diff views, inline reviews, automated CI/CD checks, and merge strategies.
@@ 2.4. Mining GitHub Issues and Pull Requests
GitHub repositories provide a rich data source for understanding developer challenges, collaboration and software evolution. Prior work has mined issues and PRs to study collaboration patterns (Tsay et al., 2014), pull-based development models, and software evolution dynamics (Gousios et al., 2014). Domain-specific studies have examined repositories for identifying deverlopers pain points for deep learning frameworks (Han et al., 2020), Quantum Software Engineering (Li et al., 2021), AI systems (Yang et al., 2023), and scientific workflow systems (Alam et al., 2025). These works demonstrate how textual and socio-technical mining can reveal developer and user needs. Our study contributes to this literature by systematically analyzing over 25K issues and PRs from nf-core, providing the first large-scale taxonomy of user concerns in pipeline repositories.
@@ 2.5. Topic Modeling
Topic modeling is a statistical technique for finding latent themes in large text collections, facilitating organization and understanding of unstructured content (Jelodar et al., 2019). It has been widely applied in document clustering, information retrieval, recommendation systems, and software engineering research, particularly in mining software repositories and source code analysis (Asuncion et al., 2010; Bagheri et al., 2014; Gethers and Poshyvanyk, 2010). Traditional models such as Latent Dirichlet Allocation (LDA) and its variants (e.g., sLDA, CTM, RTM, DTM) (Blei et al., 2003) have been extensively used but suffer from fixed topic numbers, lack of contextual understanding, and limited scalability. They are prone to generating incoherent topics, particularly in the presence of noisy or extensive datasets, and their performance is susceptible to hyperparameter settings, requiring extensive experimentation (Alam et al., 2025). In contrast, BERTopic (Grootendorst, 2022) leverages transformer-based embeddings, and produces more coherent, context-aware, and interpretable topics. Its ability to handle technical, short, and noisy texts such as GitHub data makes it particularly suitable for this study.
@@ 2.6. Topic Analysis of GitHub Data
Topic modeling with GitHub data involves analyzing text from repositories, issues, PRs, and comments to identify key topics. This technique can uncover trends, common issues, or popular features within software development communities. Researchers apply algorithms like LDA, BERTopic, and others to categorize the vast amount of unstructured text data on GitHub to understand challenges in various domains, such as quantum software engineering (Li et al., 2021), cryptography APIs (Nadi et al., 2016), open-source AI repositories (Yang et al., 2023), deep learning frameworks (Han et al., 2020), desktop web apps (Scoccia et al., 2021), scientific workflow systems development (Alam et al., 2025) and others (Dhasade et al., 2020; Jokhio, 2021; Wang et al., 2019). In addition, the mining of GitHub repositories has become a cornerstone in empirical research for uncovering collaboration patterns (Gousios et al., 2014), studying software evolution (Kalliamvakou et al., 2014), and characterizing issue and PR management practices (Tsay et al., 2014).
3. Methodology
@@ 3.1. Data Collection
As of August 22, 2025, the nf-core framework comprises 137 pipelines, of which 12 have been archived. Since archived pipelines are no longer actively maintained, they are excluded from our analysis, resulting in a final dataset of 125 active pipelines. Using the GitHub REST API, we collect a total of 25,586 issues and PRs across these pipelines. However, as many PRs are directly linked to existing issues, treating them as independent records could introduce redundancy. To address this, we employ the GitHub Search API to identify and exclude 410 PRs that are explicitly associated with existing issues. Thus, we have 25,176 entries for the further analysis. Among all pipelines, nf-core/sarek (Hanssen et al., 2024) has the highest number of issues and PRs (1,947), followed by nf-core/rnaseq (Ewels et al., 2020b) with 1,583. To enrich our dataset with repository-level context, we systematically retrieve metadata and issue/PR information for each repository under study. The collected metadata includes key attributes such as the number of contributors, stars, forks, as well as watchers of each repository. It is important to note that nf-core pipelines are continuously evolving. For this study, we collect issues and PRs associated with nf-core pipelines spanning from March 2018, coinciding with the establishment of nf-core, through August 22, 2025. We share the data collection scripts in the replication package (Random, 2025) for transparency and reproducibility.

### v2

@@ Abstract.
Scientific Workflow Systems (SWSs) such as Nextflow have become essential software frameworks for conducting reproducible, scalable, and portable computational analyses in data-intensive fields like genomics, transcriptomics, and proteomics. Building on Nextflow, the nf-core community curates standardized, peer-reviewed pipelines that follow strict testing, documentation, and governance guidelines. Despite its widespread adoption, little is known about the challenges users face in developing and maintaining these pipelines. This paper presents an empirical study of 25,173 issues and pull requests from these pipelines to uncover recurring challenges, management practices, and perceived difficulties. Using BERTopic modeling, we identify 13 key challenges, including pipeline development and integration, bug fixing, integrating genomic data, managing CI configurations, and handling version updates. We then examine issue-resolution dynamics, showing that 89.38% of issues and pull requests are eventually closed, with half resolved within 3 days. Statistical analysis reveals that the presence of labels (large effect, $\mathit{d}=0.94$) and code snippets (medium effect, $\mathit{d}=0.50$) significantly improves the likelihood of resolution. Further analysis reveals that tool development and repository maintenance poses the most significant challenges, followed by testing pipelines and CI configurations, and debugging containerized pipelines. Overall, this study provides actionable insights into the collaborative development and maintenance of nf-core pipelines, highlighting opportunities to enhance their usability, sustainability, and reproducibility.
@@
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793376isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Scientific Workflow Management Systemsccs: Applied computing Bioinformatics Pipelines
@@ 1. Introduction
The rapid growth of high-throughput sequencing and other omics (e.g., genomics, transcriptomics, and metagenomics) technologies has led to an explosion of large, heterogeneous, and complex bioinformatics datasets (Yamada et al., 2021). Analyzing these data requires chaining together multiple computational tools (modules) with diverse dependencies, formats, and parameter configurations. To address these challenges, scientific workflow systems (SWSs) such as Nextflow (Di Tommaso et al., 2017), Snakemake (Köster and Rahmann, 2012), and Galaxy (54) have emerged as indispensable software frameworks, enabling researchers to design, execute, and share multi-step computational analyses with improved reproducibility, scalability, and portability (Deelman et al., 2009; Ewels et al., 2020a).
One of the most widely adopted SWSs is Nextflow (Di Tommaso et al., 2017), which provides a domain-specific language for defining pipelines and integrates seamlessly with package managers (e.g., Conda) and containerization technologies (e.g., Docker, Singularity). Nextflow also supports deployment on various infrastructures, including local workstations, high-performance computing (HPC) clusters, and cloud platforms. This flexibility allows developers to prototype pipelines locally and subsequently scale them to large computational resources with minimal configuration changes.
@@
From a development standpoint, nf-core offers a standardized pipeline template that enforces coding style and guidelines, reducing entry barriers and promoting code consistency. Automated synchronization propagates template updates, while governance, code reviews, and integration with Bioconda/conda-forge support collaborative and sustainable development. Despite these strengths, maintaining nf-core pipelines remains challenging. Contributors must manage dependencies across diverse environments, resolve synchronization conflicts, update documentation, and ensure rigorous testing. These tasks are coordinated primarily through GitHub issues and pull requests, which facilitate bug reports, feature requests, and discussions. While prior studies have examined collaboration and maintenance through GitHub artifacts (Alam and Roy, 2025; Yang et al., 2023; Li et al., 2021; Alam et al., 2025), systematic analyses of the types of questions and issues raised within nf-core’s bioinformatics workflows are still lacking.
@@
In this paper, we address this gap through an empirical study of GitHub issues and pull requests from nf-core pipeline repositories. Using BERTopic modeling, we identify 13 key challenge areas, including pipeline development and integration, bug fixing, genome data integration, CI management, and version updates, and examine how developers manage and resolve them. Our analysis shows that 89.38% of issues and pull requests are resolved, with 50.74% closed within three days. Across 125 active repositories (12 are archived out of 137 total as of August 22, 2025), three have no open issues or pull requests, and two have none closed. About 4.64% of issues and pull requests remain entirely ignored. We also observe that 92.74% (115) of repositories utilize labeling, while 90.32% (112) assign issues and pull requests to specific contributors. Beyond the default labels, we find three additional frequently used labels among the top ten, namely WIP, DSL2, and docs. Among closed cases, 60.88% are self-closed, whereas only 17.13% involve assigned contributors. Furthermore, our analysis indicates that issues and pull requests with labels and code blocks are more likely to be closed, suggesting that providing appropriate labels and essential code snippets can enhance the clarity and resolution of reported issues and pull requests. We make the following contributions in this paper:
@@
1. We conduct an empirical study to systematically investigate issues and pull requests in open-source analysis pipelines within the nf-core ecosystem. By analyzing 25,173 issues and pull requests from 125 nf-core pipeline repositories, our study provides the workflow community with a comprehensive dataset that facilitates a deeper understanding of collaboration, maintenance, and development practices.
@@
2. We categorize user-reported challenges into a taxonomy of 13 topics, offering insights into recurring problems and development practices. This categorization aims to support developers in better understanding user challenges, thereby guiding improvements in pipeline usability, functionality, and overall maintenance.
@@
3. Our findings highlight the value of good practices in managing issues and pull requests. We recommend actively using GitHub features (e.g., labeling, assigning) for organization and accountability, and encourage contributors to provide clear, well-formatted descriptions with code snippets to improve clarity and resolution efficiency.
@@ 2. Background and Related Work
2.1. Scientific Workflow Systems and Pipelines
An SWS is a specialized software platform that enables researchers to design, execute, and manage complex computational experiments often referred to as scientific workflows (Liu et al., 2015). These systems automate repetitive tasks, integrate diverse tools, and ensure reproducibility by tracking data provenance. Examples include Galaxy (54), Nextflow (Di Tommaso et al., 2017), Snakemake (Köster and Rahmann, 2012), Pegasus (Deelman et al., 2015), and Taverna (Oinn et al., 2004).
A pipeline is a concrete instance of a scientific workflow (Barker and Van Hemert, 2007), representing a sequence of data processing steps tailored to a specific research task, such as RNA-seq analysis in bioinformatics. Pipelines can be graph-based or script-based, and developers can use domain-specific language (DSL) (Fowler, 2010), common workflow language (CWL) (Amstutz et al., 2016), workflow description language (WDL), or yet another workflow language (YAWL) (Van Der Aalst and Ter Hofstede, 2005) to develop pipelines. Community repositories such as nf-core curate standardized pipelines. Examples include rnaseq (Ewels et al., 2020b), sarek (Hanssen et al., 2024) and methylseq (nf-core methylseq authors, 2024). In short, the SWS is the platform, while the pipeline is the recipe. For example, Nextflow provides an engine for running an RNA-seq pipeline, ensuring scalability, reproducibility, and efficient execution.
@@ 2.1. Scientific Workflow Systems and Pipelines
The nf-core project develops reproducible pipelines using Nextflow (Ewels et al., 2020a). Supported by many leading organizations such as Seqera, SciLifeLab, the Center for Genomic Regulation, AWS, and Microsoft Azure, it ensures high-quality, portable analyses across diverse environments. Through shared maintenance, peer review, and centralized management, it fosters collaboration and enhances reuse. As of August 22, 2025, it hosts 125 active pipelines and over 1,600 modules (nf-core community, 2025a), exemplifying a sustainable, community-led model for bioinformatics research (nf-core community, 2025b).
@@ 2.3. Tracking Issues and Pull Requests in GitHub
Software systems are not always reliable (Gundersen et al., 2018), and maintaining open-source repositories presents ongoing challenges. To facilitate collaboration, GitHub introduced Pull Requests in 2008 (GitHub Marketplace, 2025) to propose, review, and merge code changes and Issues in 2009 (Preston-Werner, 2009) for reporting bugs, discussing features, and asking questions. Figure 1 illustrates an issue from the nf-core/mag [^2] https://github.com/nf-core/mag/issues/378 pipeline designed for the assembly, binning, and annotation of metagenomes. In this example, the individual who created the issue, referred to as the issue raiser, provides a title and a description that specify the reported concern. Any GitHub user with access to the repository may contribute to the discussion through comments, which often provide clarifications, additional details, or potential solutions. For management purposes, issues are supported by two fundamental mechanisms: labels and assignees. Labels can be created by contributors with write access to the repository and applied or removed by those with triage permissions, thereby enabling classification and prioritization. Assignees can be designated by users with write access, ensuring that responsibility for resolution is explicitly assigned. All actors involved in the lifecycle of an issue, including the raiser, commenters, and assignees, are collectively regarded as participants. Both issues and pull requests share features such as titles, descriptions, labels, and threaded discussions. Pull requests, however, add code-specific functionality, commit and diff views, inline reviews, automated CI/CD checks, and merge strategies.
@@ 2.4. Mining GitHub Issues and Pull Requests
GitHub repositories provide a rich data source for understanding developer challenges, collaboration, and software evolution. Prior work has mined issues and pull requests to study collaboration patterns (Tsay et al., 2014), pull-based development models, and software evolution dynamics (Gousios et al., 2014). Domain-specific studies have examined repositories for identifying developers’ pain points for deep learning frameworks (Han et al., 2020), Quantum Software Engineering (Li et al., 2021), AI systems (Yang et al., 2023), and scientific workflow systems (Alam et al., 2025). Our study contributes to this literature by systematically analyzing over 25K issues and pull requests from nf-core, providing the first large-scale taxonomy of user concerns in pipeline repositories.
@@ 2.5. Topic Modeling
Topic modeling is a statistical technique for finding latent themes in large text collections, facilitating organization and understanding of unstructured content (Jelodar et al., 2019). It has been widely applied in document clustering, information retrieval, recommendation systems, and software engineering research, particularly in mining software repositories and source code analysis (Asuncion et al., 2010; Bagheri et al., 2014; Gethers and Poshyvanyk, 2010). Traditional models such as LDA and its variants (e.g., sLDA, CTM, RTM, DTM) (Blei et al., 2003) have been widely used but suffer from fixed topic numbers, lack of contextual understanding, and limited scalability. They are prone to generating incoherent topics, particularly in the presence of noisy or extensive datasets, and their performance is susceptible to hyperparameter settings, requiring extensive experimentation (Alam et al., 2025). In contrast, BERTopic (Grootendorst, 2022) leverages transformer-based embeddings, and produces more coherent, context-aware, and interpretable topics. Its ability to handle technical, short, and noisy texts such as GitHub data makes it particularly suitable for this study.
@@ 2.6. Topic Analysis of GitHub Data
Topic modeling with GitHub data involves analyzing issues, pull requests, and comments to identify key topics. This technique can uncover trends, common issues, or popular features within software development communities. Researchers apply algorithms like LDA, BERTopic, and others to categorize the vast amount of unstructured text data on GitHub to understand challenges in various domains, such as QSE (Li et al., 2021), cryptography APIs (Nadi et al., 2016), open-source AI repositories (Yang et al., 2023), deep learning frameworks (Han et al., 2020), desktop web apps (Scoccia et al., 2021), scientific workflow systems development (Alam et al., 2025) and others (Dhasade et al., 2020; Jokhio, 2021; Wang et al., 2019). In addition, the mining of GitHub repositories has become a cornerstone in empirical research for uncovering collaboration patterns (Gousios et al., 2014), studying software evolution (Kalliamvakou et al., 2014), and characterizing issues and pull requests management practices (Tsay et al., 2014).
3. Study Design
@@ 3.1. Data Collection
As of August 22, 2025, the nf-core framework comprises 137 pipelines, of which 12 have been archived. Since archived pipelines are no longer actively maintained, they are excluded from our analysis, resulting in a final dataset of 125 active pipelines. Using the GitHub REST API, we collect 25,586 issues and pull requests across these pipelines. However, as many pull requests are directly linked to existing issues, treating them as independent records could introduce redundancy. To address this, we employ the GitHub Search API to identify and exclude 410 pull requests. Thus, in total, we have 25,176 issues and pull requests. Among all pipelines, nf-core/sarek (Hanssen et al., 2024) has the highest number of issues and pull requests (1,947), followed by nf-core/rnaseq (Ewels et al., 2020b) with 1,583. To enrich our dataset with repository-level context, we retrieve metadata for each repository under study. The collected metadata includes key attributes such as the number of contributors, stars, forks, and watchers of each repository. It is important to note that nf-core pipelines are continuously evolving. For this study, we collect issues and pull requests spanning from March 2018, coinciding with the establishment of nf-core, through August 22, 2025. We share the data collection scripts in the replication package (Random, 2025) for transparency and reproducibility.